Circular No. 6079

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

X-RAY NOVA IN SCORPIUS R. M. Wagner and R. Bertram, Ohio State University; P. S. Smith,

University of Arizona; and C. R. Shrader, Goddard Space Flight Center, NASA, report: “Optical spectra obtained at the Steward Observatory 2.3m telescope (range 440-870 nm, resolution 1.4 nm) on Sept. 1.1 UT, and at the Perkins 1.8-m telescope (range 450-570 nm, resolution 1 nm) on Sept. 2.1 and 8.1, of the optical counterpart of GRO J1655–40 exhibit dramatic variations in both the intensity and profile of emission lines arising from Hα (656.3 nm), Hβ (486.1 nm), He II (468.6 nm), and N III (464.0 nm). On both Sept. 1 and 2, the spectrum consists of sharp, unresolved, emission lines (FWHM ≤ 550 km/s) due to Hα (equivalent width 0.3 nm) and He II (equivalent width 0.7 nm). Hβ is weak or completely absent. In addition, many emission lines due to permitted O II are present in the region 460–470 nm. These emission lines and several interstellar absorption lines are superposed on a red continuum (V −R = +1.13). By Sept. 8, the spectrum is characterized by strong Balmer emission lines and weaker lines of He II and N III. The profiles of Hα and Hβ both exhibit a double-peaked structure. The FWHM and equivalent width of Hβ emission are ∼ 1600 km/s and 1.6 nm, respectively. To our knowledge, such dramatic emission-line profile and intensity variations have not been observed in any previous x-ray nova or transient. Our spectra obtained on Sept. 1–2 and 8 bracket the second outburst of hard x-ray emission observed on Sept. 6 (IAUC 6075). No anomalous emission lines, analogous to the Doppler-shifted emission lines of SS 433 (and which might coincide with the appearance of the extended relativistic radio structure; IAUC 6073), have been noted. However, our limited spectral coverage and the extreme southern declination of the object make observations from the northern hemisphere difficult. Additional optical spectroscopy from southern-hemisphere sites would be valuable.”

SATURN A. Sanchez-Lavega, Universidad del Pais Vasco, along with J. Gomez,

J. Lecacheux, F. Colas, and I. Miyazaki, reports that continuous CCD imaging of Saturn at several wavelengths from 0.4 to 1.0 μm has been performed at the Pic-du-Midi 1.05-m telescope, and in Spain and Japan with 0.4-m telescopes, showing (at least since mid-July) the presence of a white spot preceded by a dark column in Saturn’s equatorial zone — which on Sept. 10.01 UT was at longitude 65^{o} (System I), with a drift rate (relative to System I) of +10^{o}.2/day between July 28 and Sept. 10.

1994 September 16 Daniel W. E. Green